# SciCoQA: Quality Assurance for Scientific Paper–Code Alignment

### v1

@@ E.1 Results by Programming Language
Model Python Not Python Julia MATLAB YAML C/C++ CUDA R Shell Python+YAML AVG
# Discrepancies $483$ $46$ $12$ $10$ 99 55 44 33 33 11 $530$
GPT-5 $70.4$ $69.6$ $66.7$ $90.0$ $75.0$ $60.0$ $25.0$ $100.0$ $66.7$ $100.0$ $70.4$
GPT-5 Mini $65.4$ $67.4$ $75.0$ $90.0$ $66.7$ $60.0$ $0.0$ $100.0$ $33.3$ $100.0$ $65.7$
Gemini 2.5 Pro $49.3$ $39.1$ $33.3$ $60.0$ $50.0$ $40.0$ $0.0$ $33.3$ $33.3$ $0.0$ $48.3$
GPT-5 Codex $49.1$ $43.5$ $25.0$ $70.0$ $50.0$ $40.0$ $25.0$ $66.7$ $33.3$ $100.0$ $48.7$
GPT-OSS 20B $46.6$ $50.0$ $50.0$ $60.0$ $66.7$ $60.0$ $25.0$ $33.3$ $0.0$ $0.0$ $46.8$
GPT-OSS 120B $45.3$ $39.1$ $33.3$ $80.0$ $41.7$ $40.0$ $0.0$ $33.3$ $0.0$ $0.0$ $44.7$
Gemini 2.5 Flash $42.0$ $30.4$ $33.3$ $40.0$ $33.3$ $40.0$ $0.0$ $66.7$ $0.0$ $0.0$ $40.9$
GPT-5 Nano $28.8$ $30.4$ $25.0$ $30.0$ $58.3$ $40.0$ $25.0$ $0.0$ $0.0$ $0.0$ $28.9$
Nemotron Super 49B v1.5 $24.0$ $28.3$ $25.0$ $20.0$ $41.7$ $40.0$ $50.0$ $33.3$ $0.0$ $0.0$ $24.3$
Gemini 2.5 Flash Lite $25.3$ $26.1$ $16.7$ $40.0$ $50.0$ $40.0$ $0.0$ $0.0$ $0.0$ $100.0$ $25.5$
Qwen3 30B Coder $23.4$ $23.9$ $33.3$ $20.0$ $41.7$ $0.0$ $0.0$ $33.3$ $33.3$ $0.0$ $23.4$
Qwen3 30B Inst. $24.0$ $17.4$ $25.0$ $10.0$ $33.3$ $0.0$ $0.0$ $33.3$ $0.0$ $0.0$ $23.4$
Nemotron Nano 9B v2 $16.8$ $26.1$ $25.0$ $40.0$ $50.0$ $0.0$ $25.0$ $0.0$ $0.0$ $0.0$ $17.5$
DeepSeek R1 32B $14.1$ $19.6$ $8.3$ $20.0$ $25.0$ $20.0$ $0.0$ $66.7$ $0.0$ $0.0$ $14.5$
Qwen3 30B Think. $16.1$ $15.2$ $8.3$ $30.0$ $16.7$ $20.0$ $0.0$ $0.0$ $0.0$ $0.0$ $16.0$
Qwen3 4B Inst. $16.8$ $8.7$ $8.3$ $10.0$ $0.0$ $0.0$ $0.0$ $33.3$ $33.3$ $0.0$ $16.0$
Qwen3 4B Think. $11.4$ $6.5$ $0.0$ $0.0$ $41.7$ $0.0$ $0.0$ $0.0$ $0.0$ $0.0$ $10.9$
DeepSeek Coder 16B V2 $9.3$ $4.3$ $8.3$ $10.0$ $0.0$ $0.0$ $0.0$ $0.0$ $0.0$ $0.0$ $8.9$
DeepSeek R1 8B $6.8$ $2.2$ $0.0$ $10.0$ $0.0$ $0.0$ $0.0$ $0.0$ $0.0$ $0.0$ $6.4$
Devstral 24B Small $13.3$ $2.2$ $0.0$ $0.0$ $0.0$ $0.0$ $0.0$ $33.3$ $0.0$ $0.0$ $12.3$
Magistral 24B Small $11.0$ $8.7$ $8.3$ $0.0$ $25.0$ $0.0$ $0.0$ $0.0$ $0.0$ $0.0$ $10.8$
@@
Appendix F Extended Results
F.1 Real vs Synthetic Recall
@@
To assess the validity of our synthetic data generation pipeline, we analyze the relationship between model performance on real-world discrepancies versus synthetic discrepancies. Fig. 8 visualizes the recall and relationship between the subsets. We observe a high positive Pearson correlation of $r=0.94$. This indicates that the relative ranking of models remains consistent regardless of the data origin; models that demonstrate stronger reasoning capabilities on our synthetic injections are reliably better at identifying real-world discrepancies. While this trend might be influenced by the general capabilities of the models the strong correlation confirms that the synthetic data does not introduce noise or bias that alters the model rankings. Consequently, the synthetic subset serves as a reliable proxy for evaluating performance, justifying its use for scaling the benchmark to scientific domains where real-world examples are scarce.
@@
Figure 8: Correlation between model recall on the synthetic (x-axis) and real (x-axis) subsets of SciCoQA. Each point represents one of the 21 evaluated models. The dashed line visualizes the Pearson correlation.
F.2 Detailed Results per Model
@@
# Real/Synthetic 81/530 81/530 81/0 37/0 44/0 0/530 0/280 0/250 55/12 51/26 424/43
GPT-5 $4.1$ $67.1$ $45.7$ $62.2$ $31.8$ $70.4$ $65.7$ $75.6$ $70.1$ $50.6$ $69.4$
GPT-5 Mini $4.7$ $62.8$ $44.4$ $59.5$ $31.8$ $65.7$ $59.6$ $72.4$ $70.1$ $46.8$ $64.5$
Gemini 2.5 Pro $3.1$ $47.1$ $39.5$ $51.4$ $29.5$ $48.3$ $43.2$ $54.0$ $44.8$ $33.8$ $49.7$
GPT-5 Codex $2.0$ $46.2$ $29.6$ $35.1$ $25.0$ $48.7$ $44.6$ $53.2$ $62.7$ $29.9$ $46.5$
GPT-OSS 20B $5.1$ $46.2$ $42.0$ $48.6$ $36.4$ $46.8$ $41.1$ $53.2$ $61.2$ $37.7$ $45.4$
GPT-OSS 120B $4.5$ $43.5$ $35.8$ $45.9$ $27.3$ $44.7$ $41.1$ $48.8$ $49.3$ $33.8$ $44.3$
Gemini 2.5 Flash $3.5$ $40.3$ $35.8$ $48.6$ $25.0$ $40.9$ $34.3$ $48.4$ $49.3$ $29.9$ $40.7$
GPT-5 Nano $3.4$ $27.7$ $19.8$ $24.3$ $15.9$ $28.9$ $24.6$ $33.6$ $37.3$ $15.6$ $28.3$
Nemotron Super 49B v1.5 $4.6$ $23.6$ $18.5$ $21.6$ $15.9$ $24.3$ $23.9$ $24.8$ $34.3$ $11.7$ $24.0$
Gemini 2.5 Flash Lite $2.5$ $23.9$ $13.6$ $27.0$ $2.3$ $25.5$ $19.6$ $32.0$ $31.3$ $13.0$ $24.6$
Qwen3 30B Coder $3.8$ $23.1$ $21.0$ $24.3$ $18.2$ $23.4$ $22.9$ $24.0$ $37.3$ $9.1$ $23.3$
Qwen3 30B Inst. $3.3$ $22.4$ $16.0$ $21.6$ $11.4$ $23.4$ $20.0$ $27.2$ $35.8$ $11.7$ $22.3$
Nemotron Nano 9B v2 $7.1$ $16.5$ $9.9$ $10.8$ $9.1$ $17.5$ $14.6$ $20.8$ $11.9$ $18.2$ $16.9$
DeepSeek R1 32B $4.0$ $15.5$ $22.2$ $32.4$ $13.6$ $14.5$ $12.5$ $16.8$ $25.4$ $11.7$ $14.8$
Qwen3 30B Think. $1.5$ $15.5$ $12.3$ $21.6$ $4.5$ $16.0$ $16.1$ $16.0$ $23.9$ $3.9$ $16.3$
Qwen3 4B Inst. $2.6$ $14.7$ $6.2$ $13.5$ $0.0$ $16.0$ $14.6$ $17.6$ $28.4$ $10.4$ $13.5$
Qwen3 4B Think. $1.4$ $10.6$ $8.6$ $8.1$ $9.1$ $10.9$ $9.3$ $12.8$ $11.9$ $10.4$ $10.5$
DeepSeek Coder 16B V2 $3.3$ $7.7$ $0.0$ $0.0$ $0.0$ $8.9$ $9.6$ $8.0$ $16.4$ $2.6$ $7.3$
DeepSeek R1 8B $3.6$ $6.1$ $3.7$ $8.1$ $0.0$ $6.4$ $7.5$ $5.2$ $9.0$ $3.9$ $6.0$
Devstral 24B Small $4.2$ $11.9$ $9.9$ $13.5$ $6.8$ $12.3$ $10.4$ $14.4$ $22.4$ $6.5$ $11.3$
Magistral 24B Small $3.5$ $10.8$ $11.1$ $18.9$ $4.5$ $10.8$ $10.4$ $11.2$ $29.9$ $7.8$ $8.6$
@@
F.3 Code Only Ablation Results

### v2

@@ C.2 Performance by Programming Language
Model Python Not Python Julia MATLAB YAML C/C++ CUDA R Shell Py+YAML AVG
# Discrepancies 496 46 12 10 9 5 4 3 3 1 543
GPT-5 $70.4$ $65.2$ $66.7$ $90.0$ $75.0$ $60.0$ $0.0$ $100.0$ $33.3$ $100.0$ $70.0$
GPT-5 Mini $63.7$ $69.6$ $75.0$ $90.0$ $66.7$ $60.0$ $25.0$ $100.0$ $33.3$ $100.0$ $64.3$
Gemini 3.1 Pro $56.9$ $50.0$ $50.0$ $80.0$ $50.0$ $60.0$ $0.0$ $66.7$ $0.0$ $100.0$ $56.4$
GPT-OSS 20B $46.8$ $52.2$ $50.0$ $60.0$ $66.7$ $60.0$ $50.0$ $33.3$ $0.0$ $0.0$ $47.1$
Gemini 2.5 Pro $48.8$ $45.7$ $41.7$ $60.0$ $50.0$ $40.0$ $50.0$ $33.3$ $33.3$ $0.0$ $48.4$
GPT-5 Codex $48.8$ $45.7$ $33.3$ $70.0$ $50.0$ $40.0$ $25.0$ $66.7$ $33.3$ $100.0$ $48.6$
GPT-OSS 120B $44.6$ $39.1$ $33.3$ $80.0$ $41.7$ $40.0$ $0.0$ $33.3$ $0.0$ $0.0$ $44.0$
Gemini 2.5 Flash $42.3$ $32.6$ $33.3$ $40.0$ $33.3$ $40.0$ $0.0$ $66.7$ $33.3$ $0.0$ $41.4$
GPT-5 Nano $27.2$ $30.4$ $33.3$ $30.0$ $41.7$ $40.0$ $0.0$ $33.3$ $0.0$ $0.0$ $27.4$
Nemotron Super 49B v1.5 $24.0$ $23.9$ $16.7$ $20.0$ $16.7$ $40.0$ $50.0$ $33.3$ $33.3$ $0.0$ $23.9$
Qwen3 30B Coder $24.0$ $21.7$ $25.0$ $30.0$ $41.7$ $0.0$ $0.0$ $33.3$ $0.0$ $0.0$ $23.8$
Gemini 2.5 Flash Lite $24.0$ $19.6$ $16.7$ $20.0$ $50.0$ $20.0$ $0.0$ $0.0$ $0.0$ $100.0$ $23.8$
Nemotron Nano 9B v2 $16.1$ $23.9$ $25.0$ $40.0$ $50.0$ $0.0$ $0.0$ $0.0$ $0.0$ $0.0$ $16.8$
Qwen3 30B Inst. $22.0$ $19.6$ $25.0$ $20.0$ $33.3$ $0.0$ $0.0$ $33.3$ $0.0$ $0.0$ $21.7$
DeepSeek R1 32B $16.7$ $13.0$ $8.3$ $20.0$ $8.3$ $0.0$ $0.0$ $66.7$ $0.0$ $0.0$ $16.4$
DeepSeek Coder 16B V2 $10.1$ $4.3$ $8.3$ $10.0$ $0.0$ $0.0$ $0.0$ $0.0$ $0.0$ $0.0$ $9.6$
Qwen3 4B Inst. $16.7$ $8.7$ $16.7$ $10.0$ $0.0$ $0.0$ $0.0$ $33.3$ $0.0$ $0.0$ $16.0$
Magistral 24B Small $14.5$ $10.9$ $25.0$ $0.0$ $25.0$ $0.0$ $0.0$ $0.0$ $0.0$ $0.0$ $14.2$
Devstral 24B Small $15.1$ $0.0$ $0.0$ $0.0$ $0.0$ $0.0$ $0.0$ $0.0$ $0.0$ $100.0$ $14.0$
Qwen3 30B Think. $15.5$ $15.2$ $8.3$ $30.0$ $8.3$ $20.0$ $0.0$ $33.3$ $0.0$ $0.0$ $15.5$
Qwen3 4B Think. $11.3$ $8.7$ $8.3$ $0.0$ $41.7$ $0.0$ $0.0$ $0.0$ $0.0$ $0.0$ $11.0$
DeepSeek R1 8B $5.6$ $0.0$ $0.0$ $0.0$ $0.0$ $0.0$ $0.0$ $0.0$ $0.0$ $0.0$ $5.2$
@@
Code File Prompt Position # Samples
$\leq 32k$ $253$ ($39.8\%$)
$\leq 64k$ $463$ ($72.9\%$)
$\leq 131k$ $568$ ($89.4\%$)
$\leq 262k$ $631$ ($99.4\%$)
$\leq 524k$ $635$ ($100.0\%$)
Table 5: Number of data samples by position of the last relevant code file in the prompt.
C.3 Truncation Analysis
In §5 we analyzed the model performance by input length, finding that longer inputs degrade performance. We extend this analysis by investigating the performance by the position of the last relevant code file in the prompt. The relevant code files are obtained in the validation/rephrasing step for the real data, and during discrepancy generation for the synthetic data. Table 5 shows the distribution of the position of the last relevant code file in the prompt. We find that $89.4\%$ of samples in our data contain all relevant code files within a context window of $131k$ tokens, and $99.4\%$ within a $262k$ context window.
We further analyze the recall by the position of the last relevant code file in the prompt and bin the samples and report results in Table 8. We observe substantial recall degradation even for models that experience virtually zero truncation. GPT-5 drops from $65.8\%$ to $50.7$ and Gemini 3.1 Pro from $55.0\%$ to $31.3\%$ as the relevant file appears later in the prompt.
Both the code file position distribution as well as the performance degradation demonstrate that the dominant factor is long-context reasoning, not truncation, consistent with our analysis in Fig. 4 and prior findings on long-context degradation (e.g. Liu et al. (2024b); Hsieh et al. (2024); Levy et al. (2024)).
C.4 Real vs Synthetic Recall
@@
To assess the validity of our synthetic data generation pipeline, we analyze the relationship between model performance on real-world discrepancies versus synthetic discrepancies. Fig. 9 visualizes the recall and relationship between the subsets. We observe a high positive Pearson correlation of $r=0.94$. When excluding the GPT-5 family (i.e. GPT-5, GPT-5 Mini, GPT-5 Nano, and GPT-5 Codex) from the correlation analysis, the Pearson correlation rises to $r=0.98$. This indicates that the relative ranking of models remains consistent regardless of the data origin; models that demonstrate strong capabilities on our synthetic injections are reliably better at identifying real-world discrepancies. While this trend might be influenced by the general capabilities of the models the strong correlation confirms that the synthetic data does not introduce noise or bias that alters the model rankings. Consequently, the synthetic subset serves as a reliable proxy for evaluating performance, justifying its use for scaling the benchmark to scientific domains where real-world examples are scarce.
@@
Figure 9: Correlation between model recall on the synthetic (x-axis) and real (y-axis) subsets of SciCoQA. Each point represents one of the $22$ evaluated models. The dashed lines show linear fits across all models (blue) or excluding the GPT-5 family (orange). The $r$ values denote Pearson correlation coefficients.
C.5 Gemini vs GPT Ground Truth Recall
A potential concern with the LLM-rephrased ground truth discrepancy descriptions is that a model from the same family might have an advantage when evaluated against the ground truth (GT) that was generated by the same model or from the same family. To test this, we compare per-model recall when evaluating against only the GPT-generated versus Gemini-generated ground truth descriptions across the $80$ discrepancies for which both are available (Table 7). We find no evidence of systematic provider bias. In fact, the closed OpenAI models perform the same or better against the Gemini-generated ground truth. Gemini 3.1 Pro performs slightly better against its own generations ($\Delta=-1.3$), but the other Gemini models do not exhibit any bias towards these descriptions. Among third-party model families, which serve as unbiased controls since they are not affiliated with either GT provider, results are similarly balanced: DeepSeek and NVIDIA models show no consistent preference, while the Qwen models slightly favor GPT GT (mean $\Delta=+2.0$pp). Across all $22$ models, the overall mean $\Delta$ is +0.2+0.2pp with a maximum absolute difference of just 3.73.7pp, and $9$ models favor GPT GT versus $8$ favoring Gemini GT ($5$ tied). These small, inconsistent differences suggest that the choice of ground truth source does not introduce meaningful evaluation bias.
C.6 Pooled Ground Truth Results
In §5.2 we have created a pooled ground truth for $20$ NLP and CV papers by manually verifying the predicted discrepancies from GPT-5, Gemini 2.5 Pro and GPT-OSS 20B. Including the existing discrepancies from the real-world data in SciCoQA on these papers, this resulted in a total of $129$ verified paper-code discrepancies. Following the pooling methodology from information retrieval (Buckley and Voorhees, 2004), where relevance judgments are aggregated from multiple systems to approximate a more complete ground truth, we evaluate all models against the pooled set. In case multiple ground truths exist for the same discrepancy, e.g. when both GPT-5 and Gemini 2.5 Pro surfaced the same unannotated discrepancy, we consider it detected if the prediction matches at least one description. Table 6 reports the results.
Model TP FP P R F1
GPT-5 $69$ $13$ $84.1$ $63.6$ $72.4$
GPT-5 Mini $70$ $29$ $70.7$ $62.8$ $66.5$
GPT-OSS 20B $71$ $32$ $68.9$ $63.6$ $66.1$
Gemini 2.5 Pro $54$ 99 $85.7$ $48.8$ $62.2$
Gemini 3.1 Pro $54$ $17$ $76.1$ $50.4$ $60.6$
GPT-OSS 120B $55$ $35$ $61.1$ $52.7$ $56.6$
Gemini 2.5 Flash $43$ $25$ $63.2$ $41.1$ $49.8$
GPT-5 Codex $34$ 99 $79.1$ $34.9$ $48.4$
Gemini 2.5 Flash Lite $28$ 2121 $57.1$ $30.2$ $39.5$
Nemotron 49B v1.5 $32$ $25$ $56.1$ $30.2$ $39.3$
GPT-5 Nano $31$ $39$ $44.3$ $34.1$ $38.5$
Devstral 24B Small $26$ $31$ $45.6$ $27.9$ $34.6$
Qwen3 30B Inst. $26$ $35$ $42.6$ $27.1$ $33.2$
Qwen3 30B Coder $27$ $42$ $39.1$ $27.9$ $32.6$
Qwen3 30B Think. 2020 $12$ $62.5$ $18.6$ $28.7$
Qwen3 4B Inst. 2121 $33$ $38.9$ $22.5$ $28.5$
DeepSeek R1 32B 2121 $43$ $32.8$ $22.5$ $26.7$
Magistral 24B Small $25$ $62$ $28.7$ $24.0$ $26.2$
Qwen3 4B Think. $16$ $10$ $61.5$ $14.7$ $23.8$
Nemotron Nano 9B v2 $23$ $75$ $23.5$ $17.8$ $20.3$
DeepSeek Coder 16B V2 22 $16$ $11.1$ $6.2$ $8.0$
DeepSeek R1 8B 22 $12$ $14.3$ $1.6$ $2.8$
Table 6: Precision (P), Recall (R) and F1 scores on 2020 NLP and CV papers with $129$ discrepancies. Ground truths have been pooled from predictions from GPT-5, Gemini 2.5 Pro and GPT-OSS 20B. Additionally, discrepancies from the real-world data of SciCoQA have been added.
GPT-5 achieves the highest F1 ($72.4\%$), balancing precision ($84.1\%$) and recall ($63.6\%$), followed by GPT-5 Mini and GPT-OSS 20B ($66.5\%$ and $66.1\%$). As established in the pooling literature, systems contributing to the pool have a structural advantage, so the evaluations of GPT-5, Gemini 2.5 Pro, and GPT-OSS 20B should be viewed as upper bounds (Buckley and Voorhees, 2004). Nonetheless, clear patterns emerge: a precision–recall tradeoff is evident across models, with Gemini 2.5 Pro reaching the highest precision ($85.7\%$) at the cost of recall ($48.8\%$). GPT-OSS 20B and 120B are the strongest open-weight model, while the remaining open-weight models lag considerably (next best: Nemotron 49B, F1 of $39.3\%$). Despite the pool bias favoring contributing systems, recall remains the primary bottleneck: even under favorable conditions, the best model detects at most $63.6\%$ of verified discrepancies.
Model Gemini +GPT Gemini GPT $\Delta$
GPT-5 $45.0$ $45.0$ $43.8$ 1.3-1.3
GPT-5 Mini $48.8$ $47.5$ $43.8$ 3.7-3.7
Gemini 3.1 Pro $50.0$ $50.0$ $48.8$ 1.3-1.3
GPT-OSS 20B $47.5$ $45.0$ $42.5$ 2.5-2.5
Gemini 2.5 Pro $41.2$ $40.0$ $40.0$ $0.0$
GPT-5 Codex $27.5$ $27.5$ $27.5$ $0.0$
GPT-OSS 120B $42.5$ $37.5$ $40.0$ $2.5$
Gemini 2.5 Flash $36.2$ $32.5$ $33.8$ $1.3$
GPT-5 Nano $21.2$ $18.8$ $18.8$ $0.0$
Nemotron Super 49B v1.5 $26.2$ $23.8$ $21.2$ 2.5-2.5
Qwen3 30B Coder $23.8$ $17.5$ $20.0$ $2.5$
Gemini 2.5 Flash Lite $18.8$ $16.2$ $16.2$ $0.0$
Nemotron Nano 9B v2 $13.8$ $8.8$ $11.2$ $2.5$
Qwen3 30B Inst. $20.0$ $13.8$ $17.5$ $3.7$
DeepSeek R1 32B $21.2$ $17.5$ $17.5$ $0.0$
DeepSeek Coder 16B V2 $2.5$ $2.5$ $1.2$ 1.2-1.2
Qwen3 4B Inst. $10.0$ $7.5$ $8.8$ $1.2$
Magistral 24B Small $16.2$ $15.0$ $13.8$ 1.2-1.2
Devstral 24B Small $12.5$ $8.8$ $11.2$ $2.5$
Qwen3 30B Think. $13.8$ $12.5$ $13.8$ $1.3$
Qwen3 4B Think. $8.8$ $7.5$ $8.8$ $1.2$
DeepSeek R1 8B $3.8$ $3.8$ $2.5$ 1.2-1.2
Table 7: Model performance on the real-world SciCoQA data for the $80$ instances where both a GPT-5 and Gemini 3.1 Pro ground truth discrepancy description is available. The columns show the recall performance when assessed against both (Gemini+GPT), Gemini only or GPT only ground truth. A negative $\Delta$ shows preference for Gemini’s ground truth descriptions, while a positive $\Delta$ shows a preference for GPT’s.
Bin All 0-$32k$ 3232-$64k$ 6464-$131k$ $\geq 131k$
Model # Samples $635$ $243$ $202$ $118$ $72$
GPT-5 $65.8$ $77.5$ $62.9$ $53.3$ $50.7$
GPT-5 Mini $61.7$ $75.1$ $59.5$ $44.8$ $44.8$
Gemini 3.1 Pro $55.0$ $64.4$ $55.7$ $41.9$ $37.3$
GPT-OSS 20B $46.5$ $60.1$ $41.9$ $35.2$ $26.9$
Gemini 2.5 Pro $47.1$ $58.1$ $43.8$ $37.1$ $31.3$
GPT-5 Codex $45.5$ $57.7$ $41.9$ $36.2$ $25.4$
GPT-OSS 120B $43.6$ $59.3$ $38.6$ $28.6$ $23.9$
Gemini 2.5 Flash $40.5$ $51.4$ $40.5$ $24.8$ $23.9$
GPT-5 Nano $26.3$ $34.8$ $22.4$ $21.9$ $13.4$
Nemotron Super 49B v1.5 $23.9$ $29.6$ $20.5$ $20.0$ $19.4$
Qwen3 30B Coder $23.5$ $27.3$ $20.0$ $24.8$ $17.9$
Gemini 2.5 Flash Lite $22.8$ $33.6$ $18.1$ $14.3$ $10.4$
Nemotron Nano 9B v2 $16.2$ $21.7$ $13.8$ $14.3$ $6.0$
Qwen3 30B Inst. $21.3$ $26.5$ $20.5$ $12.4$ $17.9$
DeepSeek R1 32B $16.9$ $25.7$ $16.7$ $4.8$ $3.0$
DeepSeek Coder 16B V2 $8.7$ $13.8$ $4.3$ $5.7$ $7.5$
Qwen3 4B Inst. $15.1$ $21.7$ $12.9$ $8.6$ $7.5$
Magistral 24B Small $14.5$ $20.9$ $14.3$ $6.7$ $3.0$
Devstral 24B Small $13.7$ $19.0$ $11.9$ $5.7$ $11.9$
Qwen3 30B Think. $15.0$ $19.4$ $14.3$ $7.6$ $11.9$
Qwen3 4B Think. $10.6$ $15.4$ $11.4$ $3.8$ $0.0$
DeepSeek R1 8B $4.9$ $5.5$ $4.8$ $2.9$ $6.0$
Table 8: Model performance by position of the last relevant code file in the prompt. The All column considers the full SciCoQA dataset, the following columns partition the data based on the position of the last relevant code file in the prompt.
C.7 Detailed Results per Model
@@
# Real | # Synthetic 92 | 543 92 | 543 92 | 0 42 | 0 50 | 0 0 | 543 0 | 286 0 | 257 12 | 55 29 | 64 51 | 424
GPT-5 $4.6$ $65.8$ $41.3$ $54.8$ $30.0$ $70.0$ $65.7$ $74.7$ $70.1$ $51.6$ $68.0$
GPT-5 Mini $5.3$ $61.7$ $46.7$ $59.5$ $36.0$ $64.3$ $58.7$ $70.4$ $70.1$ $45.2$ $63.8$
Gemini 3.1 Pro $3.8$ $55.0$ $46.7$ $59.5$ $36.0$ $56.4$ $54.9$ $58.0$ $62.7$ $41.9$ $56.4$
GPT-OSS 20B $5.8$ $46.5$ $42.4$ $47.6$ $38.0$ $47.1$ $41.6$ $53.3$ $61.2$ $40.9$ $45.5$
Gemini 2.5 Pro $3.5$ $47.1$ $39.1$ $54.8$ $26.0$ $48.4$ $43.0$ $54.5$ $46.3$ $34.4$ $49.7$
GPT-5 Codex $2.2$ $45.5$ $27.2$ $31.0$ $24.0$ $48.6$ $45.5$ $52.1$ $59.7$ $28.0$ $46.9$
GPT-OSS 120B $5.0$ $43.6$ $41.3$ $50.0$ $34.0$ $44.0$ $40.6$ $47.9$ $52.2$ $40.9$ $42.9$
Gemini 2.5 Flash $4.0$ $40.5$ $34.8$ $47.6$ $24.0$ $41.4$ $36.0$ $47.5$ $47.8$ $33.3$ $40.8$
GPT-5 Nano $3.8$ $26.3$ $19.6$ $26.2$ $14.0$ $27.4$ $22.7$ $32.7$ $40.3$ $17.2$ $26.1$
Nemotron Super 49B v1.5 $5.1$ $23.9$ $23.9$ $28.6$ $20.0$ $23.9$ $23.8$ $24.1$ $37.3$ $15.1$ $23.8$
Qwen3 30B Coder $4.3$ $23.5$ $21.7$ $26.2$ $18.0$ $23.8$ $23.4$ $24.1$ $38.8$ $11.8$ $23.6$
Gemini 2.5 Flash Lite $2.8$ $22.8$ $17.4$ $31.0$ $6.0$ $23.8$ $19.2$ $28.8$ $32.8$ $11.8$ $23.6$
Nemotron Nano 9B v2 $6.7$ $16.2$ $13.0$ $14.3$ $12.0$ $16.8$ $15.0$ $18.7$ $13.4$ $16.1$ $16.6$
Qwen3 30B Inst. $3.8$ $21.3$ $18.5$ $21.4$ $16.0$ $21.7$ $17.8$ $26.1$ $32.8$ $15.1$ $20.8$
DeepSeek R1 32B $4.7$ $16.9$ $19.6$ $31.0$ $10.0$ $16.4$ $14.3$ $18.7$ $23.9$ $14.0$ $16.4$
DeepSeek Coder 16B V2 $3.5$ $8.7$ $3.3$ $2.4$ $4.0$ $9.6$ $9.1$ $10.1$ $19.4$ $3.2$ $8.2$
Qwen3 4B Inst. $2.8$ $15.1$ $9.8$ $14.3$ $6.0$ $16.0$ $12.9$ $19.5$ $22.4$ $11.8$ $14.7$
Magistral 24B Small $4.3$ $14.5$ $16.3$ $21.4$ $12.0$ $14.2$ $13.3$ $15.2$ $19.4$ $14.0$ $13.9$
Devstral 24B Small $4.8$ $13.7$ $12.0$ $16.7$ $8.0$ $14.0$ $11.2$ $17.1$ $25.4$ $15.1$ $11.8$
Qwen3 30B Think. $1.7$ $15.0$ $12.0$ $19.0$ $6.0$ $15.5$ $15.0$ $16.0$ $23.9$ $4.3$ $15.8$
Qwen3 4B Think. $1.6$ $10.6$ $7.6$ $7.1$ $8.0$ $11.0$ $9.1$ $13.2$ $11.9$ $7.5$ $10.9$
DeepSeek R1 8B $4.2$ $4.9$ $3.3$ $7.1$ $0.0$ $5.2$ $6.6$ $3.5$ $9.0$ $1.1$ $5.1$
@@
C.8 Code Only Ablation Results